2x^3+5x^2-x-22 x=-3
Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

2x^3+5x^2-x-22 x=-3
-----
2x^3 + 5x^2 -x - 19 = 0
----
I graphed the cubic and found only one Real solution at x = 1.587
The other solution must be complex numbers.
----------
Cheers,
Stan H.

Answer by jim_thompson5910(35256)

2x^3+5x^2-x-22

2(-3)^3+5(-3)^2-(-3)-22 ... plug in x = -3

2(-27)+5(9)-(-3)-22

-54+45+3-22

-9+3-22

-6-22

-28

So 2x^3+5x^2-x-22 = -28 when x = -3